# A Cognitively-Inspired Neural Architecture for Visual Abstract Reasoning Using Contrastive Perceptual and Conceptual Processing

### v1

@@ Abstract
We introduce a new neural architecture for solving visual abstract reasoning tasks inspired by human cognition, specifically by observations that human abstract reasoning often interleaves perceptual and conceptual processing as part of a flexible, iterative, and dynamic cognitive process. Inspired by this principle, our architecture models visual abstract reasoning as an iterative, self-contrasting learning process that pursues consistency between perceptual and conceptual processing of visual stimuli. We explain how this new Contrastive Perceptual-Conceptual Network (CPCNet) works using matrix reasoning problems in the style of the well-known Raven’s Progressive Matrices intelligence test. Experiments on the machine learning dataset RAVEN show that CPCNet achieves higher accuracy than all previously published models while also using the weakest inductive bias. We also point out a substantial and previously unremarked class imbalance in the original RAVEN dataset, and we propose a new variant of RAVEN—AB-RAVEN—that is more balanced in terms of abstract concepts.
@@
This document is a supplementary material for the paper A Cognitively-Inspired Neural Architecture for Visual Abstract Reasoning Using Contrastive Perceptual and Conceptual Processing. In this document, we provide more information about (1) the new variant of RAVEN—the AB-RAVEN dataset, (2) implementation details about the CPCNet architecture used in our experiments, (3) more experimental results on I-RAVEN and RAVEN-FAIR, and ablation studies of CPCNet.
@@ Introduction
    Using the RAVEN abstract reasoning dataset, we experimentally demonstrate that CPCNet is more effective than previous architectures by achieving the highest accuracy with the weakest inductive bias.
@@ Conclusion
In this paper, we investigated a cognitively-inspired paradigm for solving visual abstract reasoning tasks by leveraging the interplay between perceptual and conceptual processing. We discussed feasibility of this approach and detailed rationale, and, we designed a deep neural net architecture to simulate the interplay of processes suggested in the cognitively-inspired paradigm, i.e., the interplay and consistency between perceptual and conceptual processing. Experiments shows that our CPCNet architecture is more effective than all previous models for solving RAVEN. Moreover, with a balanced dataset (AB-Raven) in terms of abstract concepts, this architecture can even achieve a nearly perfect performance.

### v3

@@ Abstract
We introduce a new neural architecture for solving visual abstract reasoning tasks inspired by human cognition, specifically by observations that human abstract reasoning often interleaves perceptual and conceptual processing as part of a flexible, iterative, and dynamic cognitive process. Inspired by this principle, our architecture models visual abstract reasoning as an iterative, self-contrasting learning process that pursues consistency between perceptual and conceptual processing of visual stimuli. We explain how this new Contrastive Perceptual-Conceptual Network (CPCNet) works using matrix reasoning problems in the style of the well-known Raven’s Progressive Matrices intelligence test. Experiments on the machine learning datasets, the RAVEN family and PGM, show that CPCNet achieves higher accuracy than all previously published models while also using the weakest inductive bias. We also point out a substantial and previously unremarked class imbalance in the original RAVEN dataset, and we propose a new variant of RAVEN—AB-RAVEN—that is more balanced in terms of abstract concepts.
@@
This document is a supplementary material for the paper A Cognitively-Inspired Neural Architecture for Visual Abstract Reasoning Using Contrastive Perceptual and Conceptual Processing. In this document, we provide more information about (1) the new variant of RAVEN—the AB-RAVEN dataset, (2) implementation details about the CPCNet architecture used in our experiments, (3) more experimental results on I-RAVEN and RAVEN-FAIR, (4) ablation studies of CPCNet, and (5) Experiments on PGM.
@@ Introduction
    Using the abstract reasoning datasets–RAVEN, I-RAVEN, RAVEN-FAIR, and PGM, we experimentally demonstrate that CPCNet is more effective than previous architectures by achieving the highest accuracy with the weakest inductive bias.
@@
Besides the RAVEN family, we also tested CPCNet on PGM and achieved the best accuracy of 98.4%. More details can be found in the supplementary material.
@@ Conclusion
In this paper, we investigated a cognitively-inspired paradigm for solving visual abstract reasoning tasks by leveraging the interplay between perceptual and conceptual processing. We discussed feasibility of this approach and detailed rationale, and, we designed a deep neural net architecture to simulate the interplay of processes suggested in the cognitively-inspired paradigm, i.e., the interplay and consistency between perceptual and conceptual processing. Experiments shows that our CPCNet architecture is more effective than all previous models for solving RPM-like datasets. Moreover, with a balanced dataset (AB-Raven) in terms of abstract concepts, this architecture can even achieve a nearly perfect performance.